OpenStudy (anonymous):

find the prime factorization of 350

OpenStudy (anonymous):

7 x 5^2 x 2

OpenStudy (anonymous):

140

OpenStudy (anonymous):

5*5=25 7*25=175 175*2=350 Do you get it?